OpenStudy (lgbasallote):

i tried wolframming this but it didnt give me anything :/ \[\mathcal L \{ \cos (4t) u(t - \pi)\}\] i got something like \[-\frac{s}{s^2 + 16}\]or something like that

OpenStudy (turingtest):

\[\mathcal L\{f(t-c)u(t-c)\}=e^{-cs}F(s)\]but c isn't the same in both...

OpenStudy (lgbasallote):

i think this is more of the \[\mathcal L \{ f(t) u(t-c)\} = e^{-cs} \mathcal L \{f(t+c) \}\]

OpenStudy (anonymous):

yes it is of that form.

OpenStudy (turingtest):

I guess, in which case \(c=\pi\) and \[f(t+\pi)=\cos(4(t+\pi))=\cos(4t+4\pi)=\cos(4t)\]

OpenStudy (turingtest):

that's funny, it goes back where it started...

OpenStudy (lgbasallote):

yes that was my main concern.. \[\cos[4(t+ \pi)]\] i would like to know if this is equal to \[-\cos 4t\]

OpenStudy (turingtest):

no, it's just \[\cos(4t)\]because\[\cos(4t+4\pi)=\cos(4t+2(2\pi))\]and adding any multiple f \(2\pi\) to the angle will take us back where we started

OpenStudy (turingtest):

any *whole number* multiple of...

OpenStudy (turingtest):

so I think the final answer is just\[{se^{-cs}\over16+s^2}\]

OpenStudy (turingtest):

I could be wrong...

OpenStudy (lgbasallote):

well there can be only one right...

OpenStudy (anonymous):

turing test answer is correct.

OpenStudy (turingtest):

replace c with pi in my answer :/\[{se^{-\pi s}\over16+s^2}\]

OpenStudy (lgbasallote):

my teacher really loves teaching us wrong stuffs huh

OpenStudy (turingtest):

a more rigorous proof that adding a whole number multiple of 2pi leaves the argument the same

OpenStudy (anonymous):

yeah he is really loves that :P

OpenStudy (lgbasallote):

well he still gives the grades...

OpenStudy (anonymous):

usually when n=even \[\Large \cos(n \pi)=1\] when n is odd \[\Large \cos(n \pi)=-1\]